Q. Choose the correct answer.
Q. The vertex of parabola y=(-1)/2x2-x+4 is
a. (1,9/4)
b.( -1,9/4)
c. (9/2, -1)
d. (9/2, 9/2)
1
Expert's answer
2019-07-16T08:52:31-0400

The vertex of parabola "y=-\frac{1}{2}{x^2}-x+4" is at "(-1,\frac{9}{2})". Thus, there is no correct answer in the list of variants a, b, c, d.


Explanation:

Here, a ="-\frac{1}{2}"  and b = -1

So, the x -coordinate of the vertex is:

"x = -\frac{b}{2а} = -\frac{-1}{2*(-\frac{1}{2})}=-\frac{-1}{-1}=-1".

Substituting in the original equation to get the

y -coordinate, we get:

"y=-\frac{1}{2}{x^2}-x+4 = -\frac{1}{2}{(-1)^2}-(-1)+4 = -\frac{1}{2} + 1 +4 = 4.5 = -\frac{9}{2}"

Answer. The vertex of parabola "y=-\frac{1}{2}{x^2}-x+4" is at "(-1,\frac{9}{2})". Thus, there is no correct answer in the list of variants a, b, c, d.